five years ago,neeta was thrice as old as geeta.ten years later ,neeta will be twice as old as geeta.how old are neeta and geeta now?
Answer by ankor@dixie-net.com(22740)

let n = N's present age
let g = G's present age
:
write an equation and simplify it for each statement:
:
"five years ago,neeta was thrice as old as geeta."
n - 5 = 3(g-5)
n - 5 = 3g - 15
n = 3g - 15 + 5
n = 3g - 10
:
"ten years later ,neeta will be twice as old as geeta."
n + 10 = 2(g+10)
n + 10 = 2g + 20
n = 2g + 20 - 10
n = 2g + 10
Replace n with (3g-10) from the 1st statement
3g - 10 = 2g + 10
3g - 2g = 10 + 10
g = 20 is G's present
:
I'll let you find N's age